### v1

# Provably Overwhelming Transformer Models with Designed Inputs

Abstract
We develop an algorithm which, given a trained transformer model $\mathcal{M}$ as input, as well as a string of tokens $s$ of length $n_{fix}$ and an integer $n_{free}$, can generate a mathematical proof that $\mathcal{M}$ is “overwhelmed” by $s$, in time and space $\widetilde{O}(n_{fix}^{2}+n_{free}^{3})$. We say that $\mathcal{M}$ is “overwhelmed” by $s$ when the output of the model evaluated on this string plus any additional string $t$, $\mathcal{M}(s+t)$, is completely insensitive to the value of the string $t$ whenever length($t$) $\leq n_{free}$. Along the way, we prove a particularly strong worst-case form of “over-squashing” (Alon & Yahav, 2021; Barbero et al., 2024), which we use to bound the model’s behavior. Our technique uses computer-aided proofs to establish this type of operationally relevant guarantee about transformer models. We empirically test our algorithm on a single layer transformer complete with an attention head, layer-norm, MLP/ReLU layers, and RoPE positional encoding. We believe that this work is a stepping stone towards the difficult task of obtaining useful guarantees for trained transformer models.
ML Theory, Formal Guarantees, Transformers, Interpretability, Machine Learning, ICML
1 Introduction
Decoder-only transformers (Vaswani et al., 2023) have become an enormously popular paradigm in the past few years (Team, 2024; OpenAI, 2024; Li et al., 2022). Yet, our theoretical understanding of these models remains limited. Proving mathematically rigorous statements about transformers face challenges due to the high dimensionality and complexity of transformers. To circumvent these problems, current techniques either require simplifying assumption or very specific restrictions on the model/ dataset. In this work we propose a different approach focused toward producing rigorous and operationally relevant guarantees for specific, trained, transformer models. That is, we ask the question:
Can we develop an algorithm which can provably bound the behavior of a specific trained transformer model?
With this motivation in mind, we design a class of algorithms which can prove when a particular type of “over-squashing” phenomenon occurs in a trained transformer model. Over-squashing is a known phenomenon in graph neural networks (GNNs) in which the representation of distinct nodes becomes arbitrarily close as the number of layers grows (Alon & Yahav, 2021; Barbero et al., 2024). Recently, Barbero et al. (Barbero et al., 2024) have shown that over-squashing occurs in the limit for a transformer model without positional embeddings.
Over-squashing on its own is an important phenomenon. Yet, it does not directly constrain the model’s output behavior. In this work, we focus on a specific related phenomenon, which we call “overwhelming”. A model is “overwhelmed” by an input string if the model’s output is unchanged by concatenating with any new string of tokens of a fixed length.
Definition 1.1 (Overwhelm, Overwhelmed, Overwhelming).
For a given transformer model $\mathcal{M}$, string of tokens $s$ of length $n_{fix}$, a fixed final token $q$111 We find it convenient to fix the final token due to its importance in determining the output logits for the transformer architecture. We refer to the final token as the “query token.”, and integer $n_{free}$ we say that $\mathcal{M}$ is “overwhelmed” by $s$ if the output of the model evaluated on $s$ plus any additional string $t$ and fixed final token $q$, $\mathcal{M}(s+t+q)$, is the same regardless of the value of the string $t$ whenever length($t$) $\leq n_{free}$.
In this work, we provide concrete algorithms (algorithm 1, algorithm 2) which can produce a proof that a given token sequence overwhelms a given model, as stated in theorem 1.2 and theorem 1.3 respectively. Our algorithms work in two stages: first we upper bound a quantity which we call worst-case deviation (denoted by $\mathcal{W}$), which bounds the extent to which the logit weights output can vary. Second, we lower-bound a quantity which we call peak-to-peak difference which measures the difference in the maximum and second maximum logit weights output by the model. We thus get our first main result:
Theorem 1.2 (Formally restated in theorem 4.5).
If Algorithm 1 returns “Overwhelmed” when run on a transformer model $\mathcal{M}$, with a fixed input string of tokens $s$, final token $q$, and context length $n_{ctx}$ then $\mathcal{M}$ is, provably, overwhelmed by $s$.
Proving “overwhelming” for transformer models can have many interesting and useful applications. For example, discovering overwhelming strings can be used to find examples of model hallucination. If a fixed string causes the output to disregard even a small set of input tokens, the model will produce the wrong result for highly sensitive functions such as parity and finding bugs in code. Additionally, overwhelming strings can be used to “jailbreak” models by forcing the model to ignore part of the system prompt. Discovering these instances can be relevant in performing safety evaluations of models. Another application would use overwhelming strings to prove no-go results for prompt engineering: no “prompt” of a fixed length can equip a model $\mathcal{M}$ to properly compute a specific function, such as parity. Here the prompt will be the free string and the function input the fixed “overwhelming” string. Finally, we can also use the frequency of overwhelming strings to define a sort of “model complexity” for a given context length; the less overwhelming that occurs for a given context window, the more “powerful” a model is for that context window.
Next, we design algorithm 2, to achieve performance improvements in a restricted setting in which the $n_{free}$ tokens are chosen from the permutation set of a single string (as opposed to arbitrary tokens). This restricted setting is of practical interest because it often occurs with much shorter “overwhelming” strings. Furthermore, the ability of a transformer to distinguish between different permutations of the same string is relevant in many real-world applications. For example, if an LLM cannot distinguish between the two strings in fig. 1, then, the model cannot identify the variable $a$ as being defined before it is used.
# Code snippet
def my_func():
a = 1
b = a + 1
# Code snippet
def my_func():
b = a + 1
a = 1
Figure 1: Two snippets of code that are identical except for the order of the assignment statements.
Theorem 1.3 (Formally restated in theorem 5.2).
If algorithm 2 returns “Overwhelmed” when run on a transformer model $\mathcal{M}$, with a fixed input string of tokens $s$, permutation string $x$, query token $q$, and context length $n_{ctx}$ then $\mathcal{M}$ is, provably, overwhelmed by $s$ when the free string is restricted to be a permutation of $x$.
We focus on single-layer transformers in this paper. Our results can be straight forwardly extended to multi-layer models by approximating Lipschitz constants (Kim et al., 2021), though the bounds obtained for higher depth models are too loose to be useful in practice. New ideas are required to obtain tighter bounds for multi-layer transformers.
Before proving these theorems, we first provide a mathematical description of the model in section 2, as well as a metric of over-squashing called worst-case deviation, in section 3. In section 4, we provide the main algorithm of this paper and prove theorem 1.2. In section 5, we prove theorem 1.3 and introduce a linear program which is helpful in bounding extremal values in attention; we believe that this linear program may be of independent interest as well. In addition, in appendix F, we prove that, in the asymptotic limit, the model converges to a fixed output model when the fixed string is repeated many times before the free string.
Empirical Results: In section 6, we use our algorithm in practice on a trained, single-layer model. The algorithm checks for “overwhelming” when restricting the free string to an element of one of two permutation classes. We use fixed strings of varying sizes (up to 500 tokens): our fixed strings come from either: (1) the test set of our training data, (2) randomly sampled tokens, (3) a repeated string. We also test whether “overwhelming” persists across multiple steps of token generation.
Related Work
Gross et al. (Gross et al., 2024) also consider a concrete instance of a trained transformer model and produce a proof that the model can select the maximum token of a sequence with a lower bound on the probability of success. Barbero et al. (Barbero et al., 2024) prove that over-squashing occurs in the limit for a (multi-layer) model without positional embeddings. Though similar to our work, their results are not directly comparable to ours as they do not provide a concrete algorithm to check for over-squashing in a trained model. Refs. (Hahn, 2020; Hahn & Rofin, 2024) use Lipschitz constants to prove various properties of neural networks and provide an implicit algorithm to check for these properties. In recent year, there has also been a flurry of impossibility results related to transformers (Peng et al., 2024; Merrill & Sabharwal, 2023; Sanford et al., 2024), these works focus on general impossibility for the computational class of the transformer architecture.
2 Model
In this paper, we focus on a standard single-layer transformer model which uses RoPE positional encoding, parallel feed-forward residuals, and layer normalization. Our model closely follows the GPT-NeoX architecture (Black et al., 2022) except that: (1) we do not use an attention mask, and (2) we use a single layer-normalization for the MLP and the attention mechanism instead of two  [^2] Ref. (Black et al., 2022) intended to initially use a single layer-normalization for the MLP and attention mechanism, but, due to a bug, this was changed in the final implementation. We thus use the intended architecture in this paper..
We use relatively standard notation for transformer models where LN is the layer normalization function, E is the embedding function, $\mathrm{Unembed}$ is the unembedding function, Attn is the attention mechanism, AttnH is an attention head, MLP is the feed-forward network, and $\mathbb{I}$ is the identity function. Concrete details can be found in appendix B.
We also use the notation in Black et al. (Black et al., 2022) where we re-write the effect of RoPE via a rotation matrix $\Theta_{i,j}$ such that for $X\in\mathbb{O}^{{n_{ctx}}}$, we have
$\mathopen{}\mathclose{{}\left(\operatorname{RoPE}(XQ)\cdot\operatorname{RoPE}(K^{T}X^{T})}\right)[i,j]$
$=\vec{e}_{i}\cdot XQ\cdot\Theta_{i,j}\cdot K^{T}X^{T}\cdot\vec{e}_{j}^{T}.$
The model we study is formally defined as follows:
Definition 2.1 ($1$-Layer Transformer).
For a single layer model, $\mathcal{M}$,
$\mathcal{M}=\mathrm{Unembed}\circ(\texttt{Attn}+\text{MLP}+\mathbb{I})\circ\texttt{LN}\circ\text{E}$
where Attn has heads $\texttt{AttnH}_{1},\dots,\texttt{AttnH}_{H}$ for some number of attention heads $H$.
We also introduce notation related to one-hot vectors:
Definition 2.2 (One Hot Space).
Let ${d_{vocab}}$ be the size of the vocabulary. We define $\mathbb{O}\subset\mathbb{R}^{d_{vocab}}$ as the set of one-hot vectors. That is, $\mathbb{O}=\{\vec{e}_{i}\mid i\in[{d_{vocab}}]\}$ where $\vec{e}_{i}$ is the $i$-th standard basis vector in $\mathbb{R}^{{d_{vocab}}}$.
3 Worst-Case Deviation and a Custom Norm
When proving statements about our model, we will find it useful to talk about a variation on the Lipschitz constant and maximum absolute deviation that we term the “worst-case deviation.” Simply put, the worst-case deviation of a function $f:\mathcal{X}\to\mathbb{R}^{n}$ is the maximum distance between the outputs of $f$ on any two points in $\mathcal{X}$.
Before defining the worst-case deviation more formally, we will first define the Lipschitz constant of a function.
Definition 3.1 (Lipschitz Constant).
For a function $f:\mathcal{X}\to\mathbb{R}^{n}$ and norm $p$, we define the Lipschitz constant of $f$ as
$\mathrm{Lip}(f)_{p}:=\sup_{X_{1},X_{2}\in\mathcal{X}}\frac{\norm{f(X_{1})-f(X_{2})}_{p}}{\norm{X_{1}-X_{2}}_{p}}.$
Definition 3.2 (Worst-Case Deviation).
For a function $f:\mathcal{X}\to\mathbb{R}^{n}$, we define the worst-case deviation of $f$ as
$\mathcal{W}(f;\mathcal{X})_{p}:=\sup_{X_{1},X_{2}\in\mathcal{X}}\norm{f(X_{1})-f(X_{2})}_{p}$
It will be convenient to “lift” the input space into two parts and then find bounds on the worst-case deviation by optimizing over the two parts separately.
Definition 3.3 (Lifted Worst-Case Deviation).
For a function $f:\mathcal{X}\to\mathbb{R}^{n}$ where $\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}$ and $f$ is well defined on $\mathcal{Y}$ and $\mathcal{Z}$, we define the lifted worst-case deviation of $f$ as
$\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})_{p}:=\sup_{Y_{1},Y_{2}\in\mathcal{Y}}\sup_{Z_{1},Z_{2}\in\mathcal{Z}}\norm{f(Y_{1},Z_{1})-f(Y_{2},Z_{2})}_{p}$
In appendix C, we state and prove some useful properties of the worst-case deviation. Mainly, we note that $\mathcal{W}$ behaves mostly like a norm but has the nice property that “lifting” the domain is monotonic. I.e. for lifted domain $\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}$, $\mathcal{W}(f;\mathcal{X})\leq\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})$.
When discussing the worst-case deviation in this paper, we will use the $\infty$-norm for vectors. For convenience, we will define an analogue of the Frobenius norm for the $\infty$-norm.
Definition 3.4 (${F_{\infty}}$-norm).
For a matrix $M\in\mathbb{R}^{n\times m}$, we define the Frobenius-$\infty$ norm (${F_{\infty}}$-norm) as
$\norm{M}_{{F_{\infty}}}:=\max_{i\in[n]}\max_{j\in[m]}\absolutevalue{M_{ij}}.$ | .
4 Input Restrictions and Proving Overwhelming
In this section, we will provide an algorithm to decide “overwhelming.” Along the way, we develop a generalizable method of upper-bounding the worst-case deviation of a single layer of a transformer model under input restriction.
Input Restrictions and Designed Space
We use the notation from Analysis of Boolean Functions to denote restrictions on inputs to a function (O’Donnell, 2014). The restriction will fix certain tokens to be a specific value and leave the rest of the tokens free.
Definition 4.1 (Input Restriction, (O’Donnell, 2014) definition 3.18).
Let $f:\mathcal{X}^{n}\rightarrow\mathcal{Y}$ be some function and $J\subset[n]$ and ${\overline{J}}=[n]\setminus J$. Let $z\in\mathcal{X}^{\overline{J}}$. Then, we write $f_{J\mid z}:\mathcal{X}^{J}\rightarrow\mathcal{Y}$ (“the restriction of $f$ to $J$ given $z$”) as the subfunction of $f$ that is obtained by fixing the coordinates of ${\overline{J}}$ to the values in $z$. Given $y\in\mathcal{X}^{J}$ and $z\in\mathcal{X}^{\overline{J}}$, we write $f_{J\mid z}(y)$ as $f(y,z)$ even though $y$ and $z$ are not literally concatenated.
Throughout this paper, we will consider a specific input restriction where the first $n_{fix}$ tokens are restricted to a string ${s}$ and the last token is fixed to token $q$. We will denote said restriction as $f_{(n_{fix},{n_{ctx}})\mid{s},q}$ where $(n_{fix},{n_{ctx}})$ denotes the set $\{n_{fix}+1,\ldots,{n_{ctx}}-1\}$.
Next, it will be useful to define the set of all possible inputs under an input restriction.
Definition 4.2 (Designed Space).
Recall, from definition 2.2, that $\mathbb{O}$ is the set of all one-hot vectors of size ${d_{vocab}}$. We denote by $\mathbb{O}^{n}$ the set of matrices where each row is a one-hot vector of size ${d_{vocab}}$. Then, let $\mathbb{O}_{des}\subset\mathbb{O}^{n_{ctx}}$ designate the set of all possible inputs under a specific input restriction. That is,
$\mathbb{O}_{des}=\mathopen{}\mathclose{{}\left\{X\in\mathbb{R}^{{n_{ctx}}\times{d_{vocab}}}\mid X=\begin{bmatrix}\vec{e}_{{s}_{1}}\\ \vec{e}_{{s}_{2}}\\ \vdots\\ \vec{e}_{{s}_{s}}\\ Y\\ \vec{e}_{q}\end{bmatrix},Y\in\mathbb{O}_{free}}\right\}.$
where $\mathbb{O}_{free}$ is the space of free tokens.
4.1 Algorithm for deciding Overwhelming
Here we consider zero temperature sampling setting where we can define our model with sampling as
$\mathcal{M}^{final}(X)=\arg\max_{i\in[{d_{vocab}}]}\mathcal{M}(X)\cdot\vec{e}_{i}^{T}$
where $\mathcal{M}$ is defined in definition 2.1. The model simply selects the token with the highest logit weight rather than sampling from the output distribution.
We define the “peak-to-peak difference” to be the difference between the logit for the most likely token and the logit for the second most likely token for sample $X$.
Definition 4.3 (Peak-to-peak difference).
Let $X\in\mathbb{O}_{des}$ be any element from the restriction and $k=\mathcal{M}^{final}(X)$. Then, we let
$\mathrm{PTP}(\mathcal{M},X)=\min_{j\in[{d_{vocab}}],j\neq k}\mathcal{M}(X)\cdot\mathopen{}\mathclose{{}\left(\vec{e}_{k}^{T}-\vec{e}_{j}^{T}}\right).$
To prove a model is overwhelmed by a fixed input we bound the worst-case deviation by the peak-to-peak difference. The following theorem summarizes the bound our algorithm is tasked with verifying.
Theorem 4.4.
If
$\mathcal{W}(\mathcal{M};\mathbb{O}_{des})_{\infty}<\mathrm{PTP}(\mathcal{M},X)/2$
for some $X\in\mathbb{O}_{des}$, then the output of $\mathcal{M}^{final}$ is “overwhelmed” under the restriction.
Proof.
As $\mathcal{M}^{final}$ always selects the token with the maximum logit value, if the coordinate-wise deviation of the restricted model’s output never differs by more than $\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)/2$ for any $X\in\mathbb{O}_{des}$, the token with the second highest logit value will never exceed that of the token with the highest value. ∎
Ultimately, we will want to make use of the above theorem alongside algorithm 1 to prove the following theorem:
Theorem 4.5 (Input Restriction).
If
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$
$<\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)/2,$
then the output of model $\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed for all inputs in $\mathbb{O}_{des}$. Moreover, we can use algorithm 1 to produce an upper bound $W$ for $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$.
To prove the above theorem, we will:
  • break down the model into its components.
  • bound the worst-case deviation of each component as a function of the blowup and shift from layer-normalization.
  • use the triangle inequality of worst-case deviation to bound the worst-case deviation of the model prior to unembedding.
  • use the Lipschitz constant of the unembedding matrix to bound the worst-case deviation of the model.
The formal proof can also be found in section D.5.
Algorithm 1 Algorithm for deciding Overwhelming
  Input: Model $\mathcal{M}$, fixed string $s$, contenxt length ${n_{ctx}}$, query token $q$.
  Output: “Overwhelmed” or “Inconclusive”.
  Calculate $B^{\min},B^{\max},S^{\min},S^{\max}$ as in definition 4.6.
  Calculate pre-softmax extremal logit values $\ell^{\min}$ and $\ell^{\max}$ via algorithm 4.
  Using the above, calculate ${\beta}_{fix}^{\max}$ and ${\beta}_{fix}^{\min}$ as in definition 4.10 and lemma 4.11.
  Calculate upper-bound $W^{\texttt{Attn}}$ as in lemma 4.12.
  Calculate upper-bound $W^{f^{enc}}$ as in lemma 4.15.
  Calculate
$W=\mathrm{Lip}(\mathrm{Unembed})$
$\cdot\mathopen{}\mathclose{{}\left(W^{\texttt{Attn}}+\mathrm{Lip}({\text{MLP}})_{\infty}\cdot W^{f^{enc}}+W^{f^{enc}}}\right)$
as per lemma 4.14 and
  Sample $X\leftarrow\mathbb{O}_{des}$
  if $W<\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)/2$  then
     Return “Overwhelmed”
  else
     Return “Inconclusive”
  end if
4.2 Proof Overview of theorem 4.5 (Algorithm Correctness)
The proofs for all the lemmas and statements in this section are deferred to appendix D.
Breaking Down the Model
Discussing normalization as a function of expectation and variance is a bit cumbersome. So, we will first define an analagous “blowup” and “shift” for normalization.
Definition 4.6 (Blowup and Shift Sets).
For some $X\in\mathbb{O}_{des}$, Let $B_{j}(X)$ (for blowup) denote
$\frac{1}{\sqrt{\mathrm{Var}[X\text{E}\cdot e_{j}^{T}]+\epsilon}}\cdot\gamma$
and let $S_{j}(X)$ (for shift) denote
$\frac{-\mathbb{E}[X\text{E}\cdot e_{j}^{T}]}{\sqrt{\mathrm{Var}[X\text{E}\cdot e_{j}^{T}]+\epsilon}}\cdot\gamma+\beta.$
We define $\mathcal{B}_{j}=\{B_{j}(X)|X\in\mathbb{O}_{des}\}$, and $\mathcal{S}_{j}=\{S_{j}(X)|X\in\mathbb{O}_{des}\}$. Moreover, for ease of notation, we will define
=$\mathcal{B}=\{(B_{1}(X),\dots,B_{d_{emb}}(X)\mid X\in\mathbb{O}_{des}\}$
$\mathcal{S}=\{(S_{1}(X),\dots,S_{d_{emb}}(X))\mid X\in\mathbb{O}_{des}\}$
and the set $\mathcal{B}\mathcal{S}$ to equal
$\mathopen{}\mathclose{{}\left\{\bigg{(}(B_{1}(X),\dots,B_{d_{emb}}(X)),(S_{1}(X),\dots,S_{d_{emb}}(X))\bigg{)}}\right\}$
where $X\in\mathbb{O}_{des}$. Finally, let $B^{\min}_{j}=\min(\mathcal{B}_{j})$, $B^{\max}_{j}=\max(\mathcal{B}_{j})$, $S^{\min}_{j}=\min(\mathcal{S}_{j})$ and $S^{\max}_{j}=\max(\mathcal{S}_{j})$.
We will now break down the model in a way such that it is easier to analyze. For component ${\texttt{comp}}\in\{\texttt{Attn},\text{MLP},\mathbb{I}\}$, we will use $f^{{\texttt{comp}}}:\mathbb{O}^{n_{ctx}}\times\mathcal{B}\mathcal{S}\to\mathbb{R}^{{n_{ctx}}\times{d_{vocab}}}$ to denote the following function:
$f^{{\texttt{comp}}}(X,(B,S))={\texttt{comp}}\circ f^{enc}(X,(B,S))$
where
$f^{enc}(X,(B,S))=(X\cdot\text{E}\cdot\mathrm{diag}(B)+S).$
Note that $f^{enc}(X,(B(X),S(X)))$ is exactly equal to $\texttt{LN}(X\cdot\text{E})$ for layernorm function LN.
We can rewrite our model as
$\mathcal{M}=\mathrm{Unembed}\circ(f^{\texttt{Attn}}+f^{\text{MLP}}+f^{\mathbb{I}})$
and so, by the monotonicity of lifting for worst-case deviation,
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})$
$\leq$ $\sum_{{\texttt{comp}}}\mathcal{W}(\mathrm{Unembed}\cdot\vec{e}_{n_{ctx}}\cdot f^{{\texttt{comp}}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})$
for ${\texttt{comp}}\in\{\texttt{AttnH},\text{MLP},\mathbb{I}\}$.
4.2.1 Bounding Blowup and Shift
We can get rather straightforward bounds on the blowup and shift of the model.
Lemma 4.7 (Blowup and Shift Bounds).
We bound blowup and shift: for every $B_{j}\in\mathcal{B}_{j}$ and $S_{j}\in\mathcal{S}_{j}$
$\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\max}+\epsilon}}\leq B_{j}\leq\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\min}+\epsilon}}$
and
$\min\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)\leq S_{j}$
$\leq\max\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)$
where $\mu^{\min}_{j},\mu^{\max}_{j}$ are the lower and upper bounds on the expectation of the $j$-th column of $X\cdot\text{E}$. $\mathrm{Var}_{j}^{\min},\mathrm{Var}_{j}^{\max}$ are similarly defined for the variance. Both are bounded in lemma D.1 within appendix D.
Moreover, we let $B^{\min}$ be a vector of the minimum values of the blowup and $B^{\max}$ be a vector of the maximum values of the blowup. Define $S^{\min}$ and $S^{\max}$ similarly for the shift.
4.2.2 Bounding Attention
For a multi-headed attention mechanism, $\texttt{Attn}(X)=[\texttt{AttnH}_{1}(X);\ldots;\texttt{AttnH}_{H}(X)]$: i.e. the output is a concatenation of the individual attention head outputs. So, the infinity norm worst-case deviation is simply bounded by the maximum worst-case deviation over the individual attention heads. We thus have the following lemma:
Lemma 4.8.
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\texttt{Attn}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq\max_{h}\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}_{h}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
The rest of this section will focus on bounding the worst-case deviation of a single attention head which will be the most challenging part of the proof.
We will show how to bound the contribution of each token to the attention weights by: (1) establishing bounds for the attention weight coming from fixed tokens to the query token, (2) establishing an upper bound for the attention weight coming from free tokens to the query token. We can then demonstrate that the attention weights are heavily biased towards the fixed tokens.
We implicitly consider rotary-type positional encodings in the attention mechanism, where the positional encodings are absorbed into the calculation of the logits. As previously mentioned, we use $\ell$ to denote the logits on the query token. It will be useful, in the proof of our results, to define and compute worst-case upper and lower bounds for $\ell$, which we define formally in definition D.4. To obtain bounds $\ell^{\min}$ and $\ell^{\max}$, we provide a simple algorithm in algorithm 4 to compute a bound on the logits:
Lemma 4.9 (Bounds on Logits).
Algorithm 4 computes the minimum and maximum values of the logits $\ell$ given point-wise bounds $B^{\min}$, $B^{\max}$, $S^{\min}$, and $S^{\max}$ as specified in lemma 4.7.
We now define the most consequential value to bound the worst-case deviation of the attention mechanism.
Definition 4.10 (Softmax Extremal Values).
Let $\alpha_{free}^{\min}=\sum_{j\in(n_{fix},{n_{ctx}})}e^{\ell_{j}^{\min}}$ and $\alpha_{free}^{\max}=\sum_{j\in(n_{fix},{n_{ctx}})}e^{\ell_{j}^{\max}}$. Similarly, let $\alpha_{fix}^{\min}=\sum_{i\in[n_{fix}]\cup\{{n_{ctx}}\}}e^{\ell_{i}^{\min}}$ and $\alpha_{fix}^{\max}=\sum_{i\in[n_{fix}]\cup\{{n_{ctx}}\}}e^{\ell_{i}^{\max}}$. Then,
${\beta}_{free}^{\min}=\frac{\alpha_{free}^{\min}}{\alpha_{fix}^{\max}+\alpha_{free}^{\min}}\quad\text{and}\quad{\beta}_{free}^{\max}=\frac{\alpha_{free}^{\max}}{\alpha_{fix}^{\min}+\alpha_{free}^{\max}}$
and
${\beta}_{fix}^{\min}=\frac{\alpha_{fix}^{\min}}{\alpha_{fix}^{\min}+\alpha_{free}^{\max}}\quad\text{and}\quad{\beta}_{fix}^{\max}=\frac{\alpha_{fix}^{\max}}{\alpha_{fix}^{\max}+\alpha_{free}^{\min}}.$
These extremal values can be used to get the bounds:
Lemma 4.11 (Minimum and Maximum after Softmax).
${\beta}_{free}^{\min}\leq\sum_{j\in(n_{fix},{n_{ctx}})}\text{softmax}(\ell_{j})\leq{\beta}_{free}^{\max}$ (1)
aswell as,
${\beta}_{fix}^{\min}\leq\sum_{j\in[n_{fix}]\cup\{{n_{ctx}}\}}\text{softmax}(\ell_{j})\leq{\beta}_{fix}^{\max}.$ (2)
In the following, $E[[n_{fix}]\cup\{{n_{ctx}}\},:]\in\mathbb{R}^{n_{fix}\times{d_{emb}}}$ denotes the matrix with the rows in $[n_{fix}]\cup\{{n_{ctx}}\}$ selected.
Lemma 4.12 (Worst-case Deviation of Attention).
Worst-case deviation of attention is bounded as follows,
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot$
$\max_{B,S}\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot\mathrm{diag}(B)+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(E\cdot\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}$
$+\norm{V}_{\infty}\cdot{\beta}_{free}^{\max}\cdot\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}\}))_{\infty}$
where $V$ is the value matrix in the attention head (see appendix B).
Building on the above, we obtain a worst-case deviation bound using point-wise upper and lower bounds on $B\in\mathcal{B}$ and $S\in\mathcal{S}$.
Corollary 4.13.
Let $B^{\min}$, $B^{\max}$, $S^{\min}$, and $S^{\max}$ be as defined at the end of Definition 4.6. Then, we have
$\mathcal{W}$ $(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq$
$2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\bigg{(}\max_{B\in[B^{\min},B^{\max}]}\|E[[n_{fix}]\cup\{{n_{ctx}}\},:]$
$\cdot\mathrm{diag}(B)\|_{{F_{\infty}}}+\max\big{(}\|{S^{\min}}\|_{\infty},\|{S^{\max}}\|_{\infty}\big{)}\bigg{)}$
$+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(EB+S)\cdot V}_{{F_{\infty}}}$
$+\norm{V}_{\infty}\cdot{\beta}_{free}^{\max}\cdot\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}))_{\infty}$
where the maximization can be calculated with a simple linear program 333 We use the notation $B\in[B^{\min},B^{\max}]$ to denote the set of vectors, $B$, where $B_{j}^{\min}\leq B_{j}\leq B_{j}^{\max}$. .
4.2.3 Bounding MLP and Identity
Because the MLP and identity components are simple feed-forward networks without any “cross-talk” between tokens, we can use simple Lipschitz bounds.
Lemma 4.14 (Worst-case Deviation of MLP and Identity).
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\text{MLP}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq\mathrm{Lip}({\text{MLP}})_{\infty}\cdot\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
and
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\mathbb{I}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$=\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
4.2.4 Bounding the Encoding Function
We upper bound the worst-case deviation for $f^{enc}$ using bounds on the “blowup and shift” in lemma 4.7:
Lemma 4.15 (Worst-case deviation of $f^{enc}$).
$\mathcal{W}$ $(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq$
$\max_{t\in[{d_{vocab}}]}$ $\max_{j\in[{d_{emb}}]}\max_{B\in[\vec{0},B^{\max}-B^{\min}]}$
$\mathopen{}\mathclose{{}\left|X[t]\cdot\text{E}\cdot\mathrm{diag}(B)\cdot\vec{e}_{j}}\right|+S_{j}^{\max}-S_{j}^{\min}$
where the inner maximum term can be computed via a simple linear program.
5 Overwhelming Under Permutation Invariance
In section 4, we provided a concrete algorithm that decides overwhelming for a fixed context size ${n_{ctx}}$. In this section, we provide a more refined algorithm which can provide a tighter bound on the worst-case deviation in the restricted setting where the free tokens are allowed to be any permutation of a fixed string. Up until this point, we have been “lifting” the domain of $\mathbb{O}_{des}$ into $\mathbb{O}_{des}\times\mathcal{B}\mathcal{S}$ to prove worst-case deviation bounds: i.e. we have been separating out the effect of layer normalization when proving a fixed string to be overwhelming.
In this section, we nullify the effect of layer normalization by defining an equivalence relation on $\mathbb{O}_{des}$ where the blowup and shift sets are the same for all elements in the equivalence class. Then, we can derive a simpler upper-bound on the worst-case deviation of the model over the equivalence class as we can treat the layer normalization as a constant factor!
Definition 5.1 (Permutation Equivalence relation).
We will define the equivalence relation $\sim$ on $\mathbb{O}_{des}$. Let $X=({s}||X_{f}||\vec{e}_{q})\in\mathbb{O}_{des}$ where $X$ are the free tokens, ${s}$ are the fixed tokens, and $\vec{e}_{q}$ is the query token. Then, we define $\sim$ such that for $X=({s}||X_{f}||\vec{e}_{q}),Y=({s}||Y_{f}||\vec{e}_{q})\in\mathbb{O}_{des}$,
$X\sim Y\iff X_{f}=\pi(Y_{f})$
where $\pi$ is a permutation of the free tokens. Note that if $X\sim Y$, then $[X\cdot\text{E}]=\mathbb{E}[Y\cdot\text{E}]$ E ] and $\mathrm{Var}[X\cdot\text{E}]=\mathrm{Var}[Y\cdot\text{E}]$. Moreover, let FreeToks be the multi-set of potential free tokens for $[X]$. I.e. $\texttt{FreeToks}=\{X_{f}[1],\dots,X_{f}[n_{free}]\}$.
To bound the worst-case deviation over an equivalence class we use a similar algorithm (algorithm 2) to algorithm 1. We still bound the worst-case deviation of attention, but now the blowup and shift sets are singletons, leading to much tighter bounds. Additionally, since blowup and shift are constant, the worst-case deviations of $f^{\text{MLP}},f^{\mathbb{I}}$ are 00, and, we can use a linear program to find the extremal softmax contributions for the free tokens.
Theorem 5.2 (Input Restriction and Permutation Invariance).
If
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};[X])_{\infty}<\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)/2$
then the output of model $\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed for all inputs in $[X]$. Moreover, algorithm 2 produces an upper bound $W$ for $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$.
A full proof of theorem 5.2 is provided in section E.2.
Algorithm 2 Algorithm for deciding Overwhelming
  Input: $\mathcal{M},{s},q$
  Output: Return “Overwhelemed” or “Inconclusive”
  Set $B=\frac{\gamma}{\sqrt{\mathrm{Var}[X\cdot E]+\epsilon}}$
  Set $S=-\mathbb{E}[X\cdot E]B+\beta$
  Set $f^{enc}(\vec{e})=\vec{e}\cdot E\cdot\mathrm{diag}(B)+S$
  Calculate ${\alpha}_{free}^{\min},{\alpha}_{free}^{\max}$ using the naive algorithm (algorithm 6) or the linear program (algorithm 3)
  Let
$\alpha_{fix}=\exp\mathopen{}\mathclose{{}\left(f^{enc}(\vec{e}_{q})\cdot Q\Theta_{{n_{ctx}},{n_{ctx}}}K^{T}f^{enc}(\vec{e}_{q})}\right)$
$+\sum_{j\in[n_{fix}]}\exp\mathopen{}\mathclose{{}\left(f^{enc}(\vec{e}_{q})\cdot Q\Theta_{i,{n_{ctx}}}K^{T}f^{enc}(\vec{e}_{{s}_{j}})}\right)$
  Calculate ${\beta}_{free}^{\max}$, ${\beta}_{fix}^{\min}$, ${\beta}_{fix}^{\min}$
  Calculate $W^{\texttt{Attn}}$ as in lemma 5.4
  Calculate $W=\mathrm{Lip}(\mathrm{Unembed})\cdot W^{\texttt{Attn}}$
  if  $W<\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)/2$  then
     Return “Overwhelmed”
  else
     Return “Inconclusive”
  end if
5.1 Bounding Attention
To bound the worst-case deviation of attention, we provide two algorithms to bound ${\alpha}_{free}^{\min},{\alpha}_{free}^{\max}$ as defined in lemma 4.11. The first, in algorithm 6, takes a “naive” approach by iterating through the position of the free tokens and finding an extremal value for each position. The second, in algorithm 3, uses a linear program to get a tighter bound. To see why algorithm 3 provides a valid bound, consider a similar program except where an integer linear program is used. Then, the integer version of algorithm 3 finds the permutation of free tokens which maximizes (resp. minimizes) the pre-softmax logits. Relaxing to a linear program, we have an upper-bound (resp. lower-bound) on the pre-softmax logits. We can formally state the correctness of algorithm 6 and algorithm 3 in the following lemma:
Lemma 5.3 (Correctness of algorithm 3 and algorithm 6).
Both algorithm 3 and algorithm 6 provide valid bounds on ${\alpha}_{free}^{\min}$ and ${\alpha}_{free}^{\max}$.
Algorithm 3 Linear Program to find ${\alpha}_{free}^{\min}$.
To find ${\alpha}_{free}^{\max}$, switch the $\min$ to a $\max$ in the objective.
  Input: $f^{enc}$ and model $\mathcal{M}$
  Return the optimal value to the following linear program:
$\min\quad$ $\sum_{j\in(n_{fix},{n_{ctx}})}$x_j, t
subject to $\forall j,\sum_{t\in\texttt{FreeToks}}x_{j,t}=1$
$\forall t,\sum_{j\in(n_{fix},{n_{ctx}})}x_{j,t}=1,$
$\forall j,t,x_{j,t}\geq 0.$
Then, with the bounds on the softmax contributions, we can state a specialized version of the bound on attention’s worst-case deviation for the permutation invariant case.
Then, we can state the following lemma:
Lemma 5.4 (Worst-case Deviation of Attention).
Worst-case deviation of attention is bounded as follows for fixed blowup $B$ and shift $S$ when considering the permutation class $[X]$ with free tokens FreeToks:
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot B+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\norm{(E[(n_{fix},{n_{ctx}}),:]B+S)\cdot V}_{{F_{\infty}}}.$
The proof follows analogously to that of lemma 4.12 except for two main differences: (1) we have a further restriction on the free tokens and thus only need to consider $E[(n_{fix},{n_{ctx}}),:]$ instead of $E$ and (2) becuase the blowup and shift are fixed, $\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S})_{\infty}=0$ and so wecan drop the last term in the original bound.
After which the proof of theorem 5.2 follows analogously to that of theorem 4.5.
6 Evaluating on a Single Layer Model
In this section, we empirically demonstrate our algorithm on a single layer transformer model trained for next token prediction on a standard text corpus. Specifically, we run algorithm 2 to calculate a bound on the worst-case deviation and peak-to-peak difference for the model where the domain has an input restriction with free tokens drawn from a single permutation class as in section 5. When algorithm 2 outputs ``$Overwhelmed^{\prime\prime}$ it follows from theorem 5.2 (informally, theorem 1.3 in the Introduction) that the output of the model is the same for all permutations.
6.1 Experimental Setup
We train a single-layer transformer model with an embedding dimension of 512 and the BERT tokenizer (Devlin et al., 2019). The model’s architecture is outlined in section 2. We train on the AG News dataset 444see PyTorch’s documentation, using the training split with a batch size of 8, learning rate of 5e-5 using the Adam optimizer, and $20$ epochs.
We examine a few different types of input restrictions 555 The first token is always fixed to the “BOS” token which connotes the beginning of a string in the BERT tokenizer. For simplicity, we think of the BOS token as part of the fixed tokens. :
  • Random String: Randomly sample alpha-numeric tokens (including space and punctuation) to form ${s}$
  • Random Sentences: Sample (and concatenate) sentences from the AG News testing set, to form ${s}$
  • Repeating Tokens: Form ${s}$ by repeating the string “what is it” (which is $3$ tokens in the BERT tokenizer).
In each case, we have a tunable number of tokens in ${s}$ [^6] Larger input restrictions are produced by concatenation.. For the permutation class, we consider the following two cases:
  • Hamlet: We use the famous quote from Shakespeare’s Hamlet: “To be or not to be, that is the question. Whether it is a nobler in the mind to suffer the slings and arrows”
  • The Old Man and the Sea: We fix the free tokens to be a snippet from the opening chapter of Hemingway’s The Old Man and the Sea: “The blotches ran well down the sides of his face and his hands had the deep-creased scars from handling heavy fish on the cords.”
Finally, the question mark token “?” is used as the query token.
In fig. 2, we show our algorithm in action for the “Hamlet” permutation class and repeated fixed string. In appendix A, we plot the worst-case deviation in fig. 3 and the peak-to-peak difference in fig. 4 for each of the input restrictions and permutation classes.
Figure 2: Worst-case deviation for the “Hamlet” permutation string and fixed repeated string. The $x$-axis is the number of total tokens and the $y$-axis is the log-scale worst-case deviation. We mark cases of provable overwhelming with a red “x.”
Observations
The experimental results for worst-case deviation bounds in fig. 3 and peak-to-peak difference in fig. 4 provide interesting insights and questions. The upper-bound on worst-case deviation trend downwards as the fixed input token size is increased. But only the “repeated-string” setting is monotonically decreasing. Moreover, all plots exhibit a “phase transition:” at around 180 tokens and again around 350 tokens, the upper-bound on worst-case deviation has a sharp drop. We note that the linear program seems to be the most useful for the “AG News” dataset when paired with the The Old Man and the Sea free string. Finally, we note that in both plots, “not much” seems to happen prior to 180 tokens. Specifically, both the peak-to-peak difference and bound on worst-case deviation remain flat for the smaller fixed string.
6.2 Overwhelming Continues Through Generation
When a transformer model is “overwhelmed”the generated token immediately after the text is fixed for any free string. However, this does not apriori imply that subsequent tokens when the model is used to repeatedly generate tokens will be fixed.
To test this we run examples of text generation where we begin with an overwhelming string $s$ and use the model to generate text. We then check if the model remains “overwhelmed”as the generated text is included as part of the fixed string. The results of these tests are plotted in fig. 5 within appendix A. One would hope the model remains overwhelmed as the fixed string is increased by the addition of the newly generated tokens. This occurs for the Hamlet free string with the repeated and random fixed strings, where as expected the worst-case deviation remains mostly constant throughout token generation. For the longer free string and the AG News fixed strings the worst-case deviation peaks after a few tokens and the model stops to be provably “overwhelmed”.
7 Conclusion and Further Work
In this work, we introduce the notion of “overwhelming” transformer models. We provide algorithms to provably check if a trained one-layer transformer is overwhelmed by some fixed input string $s$ and query token $q$. We then empirically run the algorithm on a concrete single-layer model. We obtain bounds and find examples of natural overwhelming strings for this model in the restricted permutation setting.
This work is a first step in building algorithms to give provable guarantees for practical models. Improving our existing bounds and extending the algorithms to multi-layer transformers are important problems in this direction. Moreover, we believe that our approach to proving overwhelming, the use of LP bounds, and worst-case deviation metric will be independently useful for other theoretical studies of transformer based models.
Acknowledgment
The authors are grateful for helpful discussions and feedback from Jason Gross and Peter Bajcsy. LS acknowledges funding from the NSF Graduate Research Fellowship.
Appendix A Experimental Outcomes
In this appendix we provide the plots for section 6. Firstly, fig. 3 contains the plots of worst-case deviation with and without the use of the linear program, and fig. 4 plots the peak-to-peak difference. In fig. 4 whenever our algorithm proves “overwhelming”, the point is colored red and marked with a ‘x.’
Figure 3: The worst-case deviation for the model for the three different input restrictions and two different permutation classes. The $x$-axis is the number of tokens and the $y$-axis is the base-10 logarithm of the worst-case deviation.
Figure 4: The peak-to-peak difference for the model for the three different input restrictions and two different permutation classes. Red “x”s indicate that the worst-case deviation is less than half of the peak-to-peak difference and, thus, the model output is provably invariant over the permutation class. The $x$-axis is the number of tokens and the $y$-axis is the peak-to-peak deviation.
Overwhelming Through Generation
In fig. 5 we plot the bounds on worst-case deviation and the peak-to-peak difference as the model is used to continually generate text. The newly generated tokens are continuously added to the fixed string and fed again to the model to generate the next token. The model is “overwhelmed” whenever the worst-case deviation is less than the peak-to-peak difference in the plot.
Figure 5: The peak-to-peak difference and the worst-case deviation when we continue generation through multiple tokens. The $x$-axis connotes the total number of tokens ($n_{ctx}$) throughout a generation. When the line for $PTP/2$ is above the worst-case deviation line, then our algorithm provably guarantees that the output is fixed.
Appendix B Appendix for Model Details
We provide a formal definition of each of the components in the model used in this paper.
  • softmax is the softmax function
    $\text{softmax}(\vec{\alpha})[i]=\frac{e^{\vec{\alpha[i]}}}{\sum_{i=1}^{{d_{vocab}}}e^{\vec{\alpha}[i]}}$
  • ReLU is the rectified linear unit function
    $\text{ReLU}(x)=\max(0,x)$
  • LN is the layer normalization function which for matrix $X$ performs the following column wise function for columns of $X$ (i.e. $X[:,j]$):
    $\texttt{LN}(X[i,j])=\frac{X[i,j]-\mathbb{E}[X[:,j]]}{\sqrt{\mathrm{Var}[X[:,j]]+\epsilon}}\cdot\gamma+\beta$
    where $\gamma$ and $\beta$ are learned parameters and addition and division are element-wise.
  • MLP is a one-layer feed-forward network with ReLU activation. The MLP is row wise of $X$:
    $\text{MLP}(X[i])=\text{ReLU}(X[i,:]\cdot A_{enc}+b_{enc})\cdot A_{dec}$
    Depending on the model there may be additional bias terms added after the ReLU activation. For simplicitily, we will only consider one bias term prior to the ReLU activation though the results of this paper can be easily be extended to two bias terms.
  • $\operatorname{RoPE}$ (Su et al., 2024) is the rotary position encoding which applies a rotation to key and query vectors. For input $X$ at row $i$, the rotation is applied as follows:
    $\operatorname{RoPE}(X)[i,2j]=\cos(i\theta_{j})X[i,{2j}]-\sin(i\theta_{j})X[i,2j+1]$
    and
    $\operatorname{RoPE}(X)[i,2j+1]=\sin(i\theta_{j})X[i,{2j}]+\cos(i\theta_{j})X[i,2j+1]$
    where $\theta_{j}=10000^{-2j/{d_{emb}}}$ is the frequency for dimension $j$.
  • AttnH is an attention head for matrix $X\in\mathbb{R}^{{n_{ctx}}\times{d_{emb}}}$, an attention head does the following:
    $\texttt{AttnH}(X)=\text{softmax}\mathopen{}\mathclose{{}\left(\frac{\operatorname{RoPE}(XQ)\cdot\operatorname{RoPE}(K^{T}X^{T})}{\sqrt{{d_{emb}}/H}}}\right)\cdot XV$
    where the softmax is applied column-wise. As in Ref. (Black et al., 2022), we can re-write the effect of RoPE via a rotation matrix $\Theta_{i,j}$ such that
    $\mathopen{}\mathclose{{}\left(\operatorname{RoPE}(XQ)\cdot\operatorname{RoPE}(K^{T}X^{T})}\right)[i,j]=\vec{e}_{i}\cdot XQ\cdot\Theta_{i,j}\cdot K^{T}X^{T}\cdot\vec{e}_{j}^{T}$
  • Often, we have a multi-head attention mechanism which is the concatenation of $H$ attention mechanisms along the last dimension:
    $\texttt{Attn}(X)=[\texttt{AttnH}_{1}(X);\ldots;\texttt{AttnH}_{H}(X)]$
    where $\texttt{AttnH}_{h}$ is the $h$-th attention mechanism as outlined
  • $\text{E}\in\mathbb{R}^{{d_{vocab}}\times{d_{emb}}}$ is the embedding function which maps a one hot vector in $\mathbb{R}^{d_{vocab}}$ to a vector in $\mathbb{R}^{{d_{emb}}}$.
  • $\mathrm{Unembed}\in\mathbb{R}^{{d_{emb}}\times{d_{vocab}}}$ is the unembedding function which maps a vector in $\mathbb{R}^{{d_{emb}}}$ to a one hot vector in $\mathbb{R}^{d_{vocab}}$.
Appendix C Proofs for Worst-Case Deviation
Lemma C.1 (Properties of the Worst-Case Deviation).
For any functions $f,g:\mathcal{X}\to\mathbb{R}^{n}$, norm $p$ and lift to $\mathcal{Y}\times\mathcal{Z}$, we have the following properties:
  • Triangle inequality for (lifted) $\mathcal{W}$:
    $\mathcal{W}(f+g;\mathcal{X})_{p}\leq\mathcal{W}(f:\mathcal{X})_{p}+\mathcal{W}(g:\mathcal{X})_{p}.$
  • Lifting monotonicity:
    $\mathcal{W}(f;\mathcal{X})_{p}\leq\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})_{p}.$
  • Lipschitz composition: For function $g$
    $\mathcal{W}(g\circ f;\mathcal{X})_{p}\leq\mathrm{Lip}(g)_{p}\cdot\mathcal{W}(f;\mathcal{X})_{p}.$
    As a corollary, we have that for linear operators $A$,
    $\mathcal{W}(Af;\mathcal{X})_{p}\leq\norm{A}_{p}\cdot\mathcal{W}(f;\mathcal{X})_{p}.$
    as $\mathrm{Lip}(A)_{p}=\norm{A}_{p}$.
  • $p$-norm bounds for $p,q\geq 1$ and $q>p$:
    $\mathcal{W}(f;\mathcal{X})_{q}\leq\mathcal{W}(f;\mathcal{X})_{p}$
Proof of worst-case deviation properties, lemma C.1.
We will prove each of the properties in turn.
  • Triangle inequality: We can view the maximization over $\mathcal{X}$ as occuring disjointly for $f$ and $g$: I.e.
    $\mathcal{W}(f+g;\mathcal{X})_{p}$ $\leq\sup_{X_{1},X_{2}\in\mathcal{X}}\norm{f(X_{1})+g(X_{1})-(f(X_{2})+g(X_{2}))}$
    $\leq\sup_{X_{1},X_{2},X_{1}^{\prime},X_{2}^{\prime}\in\mathcal{X}}\big{[}\norm{f(X_{1})-f(X_{2})}+\norm{g(X_{1}^{\prime})-g(X_{2}^{\prime})}\big{]}$ (by triangle inequality of norms)
    $\leq\mathcal{W}(f;\mathcal{X})_{p}+\mathcal{W}(g;\mathcal{X})_{p}$
    as desired.
  • Lifting monotonicity: The proof follows from the definition of the lifted worst-case deviation:
    $\mathcal{W}(f;\mathcal{X})_{p}$ $=\sup_{(Y_{1},Z_{1}),(Y_{2},Z_{2})\in\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}}\norm{f(Y_{1},Z_{1})-f(Y_{2},Z_{2})}_{p}$
    $\leq\sup_{Y_{1},Y_{2}\in\mathcal{Y},\;Z_{1},Z_{2}\in\mathcal{Z}}\norm{f(Y_{1},Z_{1})-f(Y_{2},Z_{2})}_{p}$
    $=\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})_{p}.$
  • Lipschitz composition: We simply have that
    $\mathcal{W}(Af;\mathcal{X})_{p}=\sup_{X_{1},X_{2}\in\mathcal{X}}\norm{Af(X_{1})-Af(X_{2})}_{p}$
    $=\sup_{X_{1},X_{2}\in\mathcal{X}}\norm{A(f(X_{1})-f(X_{2}))}_{p}$
    $\leq\sup_{X_{1}^{\prime},X_{2}^{\prime}}\frac{\norm{A(X_{1}^{\prime})-A(X_{2}^{\prime})}_{p}}{\norm{X_{1}^{\prime}-X_{2}^{\prime}}_{p}}\cdot\sup_{X_{1},X_{2}\in\mathcal{X}}\norm{f(X_{1})-f(X_{2})}_{p}$
    $=\mathrm{Lip}(A)_{p}\cdot\mathcal{W}(f;\mathcal{X})_{p}.$
  • $p$-norm bounds for $p,q\geq 1$ and $q>p$: Because we restrict $f$ to be a function which outputs vectors and $\norm{\vec{x}}_{q}\leq\norm{\vec{x}}_{p}$ for $q>p$, we have that for all $X_{1},X_{2}\in\mathcal{X}$, $\norm{f(X_{1})-f(X_{2})}_{q}\leq\norm{f(X_{1})-f(X_{2})}_{p}$. So, if there exists $X_{1},X_{2}\in\mathcal{X}$ such that $\norm{f(X_{1})-f(X_{2})}_{q}=\alpha$, then there must exist $X_{1},X_{2}\in\mathcal{X}$ such that $\norm{f(X_{1})-f(X_{2})}_{p}\geq\alpha$. Thus, the supremum over $\mathcal{X}$ for $q$ is less than or equal to the supremum over $\mathcal{X}$ for $p$.
Appendix D Proofs for Generic Framework
D.1 Proofs for Blowup and Shift Bounds
We first need to bound the variance and expectation of the columns of $\text{E}\cdot X$:
Lemma D.1 (Bounds on Expectation and Variance).
Let $X\in\mathbb{O}_{des}$ be an input sequence. Then for any column $j$, the expectation $\mu_{j}$ of the $j$-th column of $\text{E}\cdot X$ satisfies:
$\frac{1}{{n_{ctx}}}$ $\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n_{fix}}\text{E}[{s}_{i}]\vec{e}_{j}^{T}+n_{free}\min_{x\in[{d_{vocab}}]}(\text{E}[x]\vec{e}_{j}^{T})+\text{E}[q]\vec{e}_{j}^{T}}\right)\leq\mu_{j}$
$\leq\frac{1}{{n_{ctx}}}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n_{fix}}\text{E}[{s}_{i}]\vec{e}_{j}^{T}+n_{free}\max_{x\in[{d_{vocab}}]}(\text{E}[x]\vec{e}_{j}^{T})+\text{E}[q]\vec{e}_{j}^{T}}\right)$
Let $\mu_{j}^{\min}$ and $\mu_{j}^{\max}$ denote the lower and upper bounds for the $j$-th column respectively. Then, we can bound the variance of the $j$-th column of $\vec{x}\text{E}$ by
$\frac{1}{{n_{ctx}}}$ $\bigg{[}\min_{\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]}\sum_{i=1}^{s}\mathopen{}\mathclose{{}\left(\text{E}[{s}_{i}]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+n_{free}\cdot\min_{x\in[{d_{vocab}}]}\mathopen{}\mathclose{{}\left(\text{E}[x]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+\mathopen{}\mathclose{{}\left(\text{E}[q]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}\bigg{]}$
$\leq\mathrm{Var}_{j}[(\text{E}\cdot X)[:,j]]\leq$
$\frac{1}{{n_{ctx}}}$ $\bigg{[}\sum_{i=1}^{s}\max_{\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]}\mathopen{}\mathclose{{}\left(\text{E}[{s}_{i}]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+n_{free}\cdot\max_{x\in[{d_{vocab}}]}\mathopen{}\mathclose{{}\left(\text{E}[x]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+\mathopen{}\mathclose{{}\left(\text{E}[q]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}\bigg{]}.$
We define $\mathrm{Var}_{j}^{\min},\mathrm{Var}_{j}^{\max}$ to be the lower and upper bounds respectively.
Proof.
The bounds on $\mu_{j}$ follow as we minimize and maximize the contribution of each free row. The proof of for variance bounds follows similarly. ∎
We now have a simple proof for lemma 4.7 restated below:
Lemma D.2 (Blowup and Shift Bounds).
We bound blowup and shift: for every $B_{j}\in\mathcal{B}_{j}$ and $S_{j}\in\mathcal{S}_{j}$
$\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\max}+\epsilon}}\leq B_{j}\leq\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\min}+\epsilon}}$
and
$\min\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)\leq S_{j}$
$\leq\max\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)$
Proof.
Given the bounds on expectation and variance, $B^{\min},B^{\max}$ follow trivially from the definition of blowup and shift sets. $S^{\min}$ (resp. $S^{\max}$) follow from a minimization (maximization) over the possible shifts given the blowup bounds. ∎
D.2 Appendix for Bounding Worst-case Deviation of $f^{enc}$
Recall the definition of blowup and shift sets from definition 4.6. Let $B^{\min},B^{\max}$ be the vectors $(B_{1}^{\min},\dots,B_{d_{emb}}^{\min})$ and $(B_{1}^{\max},\dots,B_{d_{emb}}^{\max})$ respectively. Define $S^{\min},S^{\max}$ analogously. We restate lemma 4.15 for convenience:
Lemma D.3 (Worst-case deviation of $f^{enc}$, lemma 4.15).
Let $\vec{x}\in\mathbb{O}_{des}$. Then,
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq\max_{t\in[{d_{vocab}}]}\max_{j\in[{d_{emb}}]}\max_{B\in[\vec{0},B^{\max}-B^{\min}]}\mathopen{}\mathclose{{}\left|X[t]\cdot\text{E}\cdot\mathrm{diag}(B)\cdot\vec{e}_{j}}\right|+\mathopen{}\mathclose{{}\left|S_{j}^{\max}-S_{j}^{\min}}\right|$
where the inner maximum term can be computed via a simple linear program.
Proof.
First, note that for $j\in[{d_{emb}}]$,
$\mathcal{W}(\vec{e}_{j}^{T}\cdot(f^{enc})^{T};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $=\max_{t,B,S,B^{\prime},S^{\prime}}\mathopen{}\mathclose{{}\left\|X[t]\cdot E\cdot\mathrm{diag}(B)+S-X[t]\cdot E\cdot\mathrm{diag}(B^{\prime})-S^{\prime}}\right\|_{\infty}$
$\leq max_{t,B,S,B^{\prime},S^{\prime}}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left(X[t]\cdot E\cdot(\mathrm{diag}(B)-\mathrm{diag}(B^{\prime}))+S-S^{\prime}}\right)\cdot\vec{e}_{j}}\right|$
$\leq max_{t,B,B^{\prime}}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left(X[t]\cdot E\cdot(\mathrm{diag}(B)-\mathrm{diag}(B^{\prime}))}\right)\cdot\vec{e}_{j}}\right|+|S_{j}^{\max}-S_{j}^{\min}|.$
Note then that $\mathrm{diag}(B)_{j}-\mathrm{diag}(B^{\prime})_{j}$ is constrained by $B^{\max}_{j}-B^{\min}_{j}$. We can then maximize over $j$ to get
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq\mathcal{W}(f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{F_{\infty}}$
$\leq\max_{t,B,B^{\prime}}\max_{j}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left(X[t]\cdot E\cdot(\mathrm{diag}(B)-\mathrm{diag}(B^{\prime}))}\right)\cdot\vec{e}_{j}}\right|+|S_{j}^{\max}-S_{j}^{\min}|$
as desired. ∎
D.3 Proofs and Subalgorithms for Attention Bounds
Definition D.4 ($\ell^{min}$, $\ell^{max}$).
We use $\ell^{\min}_{i}$ to denote a worst-case lower bound on the smallest possible logit at the $i$-th position of the input to the softmax in model $\mathcal{M}$. Similarly, we use $\ell^{\max}_{i}$ to denote a worst-case upper bound on the largest possible logit at the $i$-th position of the input to the softmax in model $\mathcal{M}$.
We provide the algorithm to find the extremal values $\vec{\ell}^{\min}$ and $\vec{\ell}^{\max}$ in algorithm 4. At a high level, the algorithm computes upper and lower bounds for each position in the input sequence prior to the softmax. The bounds make use of the upper and lower bounds on the blowup and shift sets, $\mathcal{B}\mathcal{S}$.
Algorithm 4 Algorithm to find $\vec{\ell}^{\min}$ and $\vec{\ell}^{\max}$. To find the minimums and maximums over $\mathcal{B}\mathcal{S}$, we use the point-wise upper and lower bounds on the blowup and shift sets alongside the bounds obtained from algorithm 5.
  Input: $\mathcal{B}\mathcal{S},\mathcal{M}$
  for $k\in[n_{fix}]$ do
     Let
$\vec{\ell}^{\min}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\min_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{{s}_{k}}E\cdot\mathrm{diag}(B)+S)^{T}$
and
$\vec{\ell}^{\max}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\max_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{{s}_{k}}E\cdot\mathrm{diag}(B)+S)^{T}$
  end for
  for $k\in(n_{fix},{n_{ctx}})$ do
     Let
$\vec{\ell}^{\min}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\min_{t\in[{d_{vocab}}]}\min_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{t}E\cdot\mathrm{diag}(B)+S)^{T}$
and
$\vec{\ell}^{\max}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\max_{t\in[{d_{vocab}}]}\max_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{t}E\cdot\mathrm{diag}(B)+S)^{T}$
  end for
  Set the extremal values of the query token:
$\vec{\ell}^{\min}_{n_{ctx}}=\frac{1}{\sqrt{{d_{emb}}}}\min_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)^{T}$
and
$\vec{\ell}^{\max}_{n_{ctx}}=\frac{1}{\sqrt{{d_{emb}}}}\max_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)^{T}$
  Return $\vec{\ell}^{\min},\vec{\ell}^{\max}$
Algorithm 5 Simple algorithm to bound the minimum of $\vec{x}A\vec{y}^{T}$ for $X^{\min}_{i}\leq\vec{x}_{i}\leq X^{\max}_{i}$ and $Y^{\min}_{i}\leq\vec{y}_{i}\leq Y^{\max}_{i}$. To find the maximum, switch the minimum to maximum and vice versa.
  Input: $A,X^{\min},X^{\max},Y^{\min},Y^{\max}$
  $m\leftarrow 0$
  for $i=1$ to $n$ do
     for $j=1$ to $m$ do
        if $A_{ij}>0$ then
           $m\leftarrow m+A_{ij}\cdot\max\mathopen{}\mathclose{{}\left(X^{\min}_{i}Y^{\min}_{j},X^{\max}_{i}Y^{\min}_{j},X^{\min}_{i}Y^{\max}_{j},X^{\max}_{i}Y^{\max}_{j}}\right)$
        else
           $m\leftarrow m+A_{ij}\cdot\min\mathopen{}\mathclose{{}\left(X^{\min}_{i}Y^{\min}_{j},X^{\max}_{i}Y^{\min}_{j},X^{\min}_{i}Y^{\max}_{j},X^{\max}_{i}Y^{\max}_{j}}\right)$
        end if
     end for
  end for
  Return $m$
Proof sketch for lemma 4.9 (correctness of algorithm 4).
In algorithm 4, for each position $k$, we compute the minimum and maximum logit for the $k$-th position by maximizing over possible blowup and shift sets and input tokens for the free tokens. Moreover, algorithm 5 computes an upper and lower bound for the restricted bilinear form which algorithm 4 uses to compute the extremal values.
The correctness of the bilinear bound in algorithm 5 follows from a series of relaxations when writing out the explicit formula for the bilinear multiplication. To prove an upper-bound, we have that
$m$ $=\sum_{i}x_{i}A_{i,j}y_{j}$
$\leq\sum_{i}\max(A_{i,j}X_{i}^{\min}Y_{i}^{\min},A_{i,j}X_{i}^{\min}Y_{i}^{\max},A_{i,j}X_{i}^{\max}Y_{i}^{\min},A_{i,j}X_{i}^{\max}Y_{i}^{\max}).$
Then, if $A_{i,j}$ is negative, we can factor the coefficient out of the maximization and flip to a minimization. Otherwise, we can simply factor $A_{i,j}$ out of the maximization. Thus, algorithm 5 gives a valid upper-bound. The lower-bound follows analogously. ∎
We now prove lemma 4.11, restated here:
Lemma D.5 (Minimum and Maximum after Softmax).
${\beta}_{free}^{\min}\leq\sum_{j\in(n_{fix},{n_{ctx}})}\text{softmax}(\ell_{j})\leq{\beta}_{free}^{\max}$ (3)
aswell as,
${\beta}_{fix}^{\min}\leq\sum_{j\in[n_{fix}]\cup\{{n_{ctx}}\}}\text{softmax}(\ell_{j})\leq{\beta}_{fix}^{\max}.$ (4)
Proof of lemma 4.11.
We will prove the right hand side of eq. 3 and the left hand side of eq. 4 as the other follow by the same proof. Note the second inequality follows from the first inequality and the fact that the sum of the attention weights is $1$.
Now, we will show the first inequality. Clearly, smaller values of $\vec{\ell}_{i}$ results in larger value of $\text{softmax}(\vec{\ell}_{j})$. Then, let $\vec{\eta}$ be a vector where $\sum_{j\in(n_{fix},{n_{ctx}}]}\text{softmax}(\vec{\eta}_{j})>\sum_{j\in(n_{fix},{n_{ctx}}]}\text{softmax}(\vec{\ell}_{j})$. Then, we have that
00 0 $\leq\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\sum_{j}e^{\vec{\eta}_{j}}}\right)-\mathopen{}\mathclose{{}\left(\sum_{j}e^{\vec{\ell}_{j}}}\right)}\right]\mathopen{}\mathclose{{}\left(\sum_{i}e^{\vec{\ell}_{i}}}\right)$
$\Rightarrow$ $\mathopen{}\mathclose{{}\left(\sum_{j}e^{\vec{\eta}_{j}}}\right)\mathopen{}\mathclose{{}\left(\sum_{i}e^{\vec{\ell}_{i}}+\sum_{j}e^{\vec{\ell}_{j}}}\right)\leq\mathopen{}\mathclose{{}\left(\sum_{j}e^{\vec{\ell}_{j}}}\right)\mathopen{}\mathclose{{}\left(\sum_{i}e^{\vec{\eta}_{i}}+\sum_{j}e^{\vec{\eta}_{j}}}\right)$
$\Rightarrow$ $\frac{\sum_{j}e^{\vec{\eta}_{j}}}{\sum_{i}e^{\vec{\ell}_{i}}+\sum_{j}e^{\vec{\eta}_{j}}}\leq\frac{\sum_{j}e^{\vec{\ell}_{j}}}{\sum_{i}e^{\vec{\ell}_{i}}+\sum_{j}e^{\vec{\ell}_{j}}}.$
Finally, we prove the bound on worst-case deviation of attention (lemma 4.12):
Lemma D.6.
Worst-case deviation of attention is bounded as follows,
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\max_{B,S}\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot\mathrm{diag}(B)+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(E\cdot\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}+{\beta}_{free}^{\max}\norm{V}_{\infty}\cdot\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}\}))_{\infty}$
where $V$ is the value matrix in the attention head (see appendix B).
Proof of lemma 4.12.
Define $\vec{p}(X)\in\mathbb{R}^{{n_{ctx}}}$ as the probability vector for the query token post-softmax. I.e.
$\vec{p}_{j}=\text{softmax}(\vec{\ell}_{j})\text{ for }j\in[{n_{ctx}}]$
where
$\vec{\ell}=\frac{\vec{e}_{{n_{ctx}}}\cdot(Y\cdot\operatorname{RoPE}(Q)\operatorname{RoPE}(K^{T})Y^{T}}{\sqrt{{d_{emb}}}}$
for $Y=X\text{E}\cdot\mathrm{diag}(B(X))+S(X)$. Moreover, let $\vec{p}(X)_{\text{free}}=\vec{p}(X)[(n_{fix},{n_{ctx}})]$ (i.e. the values corresponding to the free tokens) and $\vec{p}(X)_{\text{fix}}=\vec{p}(X)[[n_{fix}]\cup\{{n_{ctx}}\}]$ (the values corresponding to the fixed and query token). Now, we re-write the worst-case deviation of the attention head: for $X,X^{\prime}\in\mathbb{O}_{des}$, $(B,S),(B^{\prime},S^{\prime})\in\mathcal{B}\mathcal{S}$,
$\mathcal{W}$ $(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\norm{\vec{p}(X)\cdot(X\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})\cdot(X\text{E}\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V}_{\infty}$ (by definition of an attention head)
$\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V+\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)$
$+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})$
$+S^{\prime})V\bigg{\|}+\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}.$ (by triangle inequality)
Now, we will bound the two norms in the above equation separately. First, note that ${\beta}_{free}^{\min}\leq\sum\vec{p}(X)_{\text{free}}\leq{\beta}_{free}^{\max}$ and ${\beta}_{fix}^{\min}\leq\sum\vec{p}(X)_{\text{fix}}\leq{\beta}_{fix}^{\max}$ by definition of the extremal values (definition 4.10). For the case of the free tokens,
$\max_{X,X^{\prime},B,B^{\prime},S,S^{\prime}}\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}_{\infty}$
$\leq 2\cdot\max_{X,B,S}\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}_{\infty}$ (by triangle inequality)
$\leq 2\cdot\max_{B,S}\sum_{i\in(n_{fix},{n_{ctx}})}\vec{p}(X)_{\text{free}}[i]\cdot\mathopen{}\mathclose{{}\left\|(X[i,:]\text{E}\cdot\mathrm{diag}(B)+S)V}\right\|_{\infty}$ (by triangle inequality)
$\leq 2\cdot\max_{B,S}\max_{i}{\beta}_{free}^{\max}\cdot\mathopen{}\mathclose{{}\left\|(X[i,:]\text{E}\cdot\mathrm{diag}(B)+S)V}\right\|_{\infty}$ (by definition of $\vec{p}(X)_{\text{free}}$)
where the last inequality follows from the fact that $\vec{p}(X)_{\text{free}}$ is non-negative and the sum is at most ${\beta}_{free}^{\max}$. Finally, note that $\max_{i}{\beta}_{free}^{\max}\|X[i,:]\text{E}\cdot\mathrm{diag}(B)+S)V\|\leq{\beta}_{free}^{\max}\max_{t\in[{d_{vocab}}]}\|\text{E}[t,:]\cdot(\mathrm{diag}(B)+S)V\|$ as the maximum free token contribution is at most the maximum contribution from any possible token. Finally, for the free tokens, we get a bound
$2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S,t\in[{d_{vocab}}]}\norm{(\text{E}[t]\cdot\mathrm{diag}(B)+S)V}_{\infty}=2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(\text{E}\cdot\mathrm{diag}(B)+S)V}_{F_{\infty}}.$
We now bound the fixed tokens contribution:
$\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
We start by noting that the fixed tokens must be an element of ${s}\cup\{q\}$ and are, by definition, fixed for all $X$ And so,
$\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})$
$+S^{\prime})V\bigg{\|}=\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}(\vec{p}(X)_{\text{fix}}\cdot(\text{E}[s\cup\{q\},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(\text{E}[s\cup\{q\},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
$\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$ (by the triangle inequality)
where $t_{i}$ is the $i$-th fixed token. Now, we want to re-write the last line to factor out the differing blowup and shift values.
$\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
$=\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V$
-- $\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot(\mathrm{diag}(B)+S-(\mathrm{diag}(B)+S-\mathrm{diag}(B^{\prime})-S^{\prime}))V\bigg{\|}$
$\leq\max_{X,X^{\prime},(B,S)}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$+$ $\max_{X,(B,S),(B^{\prime},S^{\prime})}\sum_{i}\vec{p}(X)_{\text{fix}}[i]\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$ (By the triangle inequality)
We now bound the two terms separately. For the second term, we have that
$\max_{X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
$\leq\max_{(B,S),(B^{\prime},S^{\prime})}\max_{i}{\beta}_{fix}^{\max}\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)-(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})\bigg{\|}\cdot\norm{V}_{\infty}$ (By definition of $\vec{p}(X^{\prime})_{\text{fix}}$)
$\leq\norm{V}_{\infty}\cdot{\beta}_{fix}^{\max}\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}).$
For the first term, we get
$\max_{(B,S),X,X^{\prime}}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$=\max_{(B,S),X,X^{\prime}}\sum_{i}\bigg{\|}\mathopen{}\mathclose{{}\left(\vec{p}(X)_{\text{fix}}[i]-\vec{p}(X^{\prime})_{\text{fix}}[i]}\right)\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$=\max_{(B,S),X,X^{\prime}}\sum_{i}\mathopen{}\mathclose{{}\left(\vec{p}(X)_{\text{fix}}[i]-\vec{p}(X^{\prime})_{\text{fix}}[i]}\right)\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$\leq\max_{(B,S),i}({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
where the last inequality follows from the fact that $\vec{p}(X)_{\text{fix}}[i]-\vec{p}(X^{\prime})_{\text{fix}}[i]$ is at most ${\beta}_{free}^{\max}-{\beta}_{free}^{\min}$. Then, by definition of the ${F_{\infty}}$ norm, we have that
$\max_{(B,S),i}({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}=({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\max_{B,S}\norm{(\text{E}[s\cup\{q\},:]\cdot\mathrm{diag}(B)+S)V}_{F_{\infty}}.$
Putting the above together, we get the desired bound:
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\max_{B,S}\norm{\text{E}[{s}\cup\{q\},:]\cdot\mathrm{diag}(B)+S}_{{F_{\infty}}}$
$+2{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(E\cdot\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}.+\norm{V}_{\infty}{\beta}_{fix}^{\max}\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}).$
D.4 Proofs for MLP Bounds
Proof of lemma 4.14.
The second statement follows from $\vec{e}_{n_{ctx}}\cdot f^{\mathbb{I}}_{n_{fix}\mid r,q}=f^{enc}$ directly. Then, the first statement follows from the Lipschitz constant of the MLP:
$\norm{\text{MLP}(\vec{e}_{n_{ctx}}\text{E}B+S)-\text{MLP}(\vec{e}_{n_{ctx}}\text{E}B^{\prime}+S^{\prime})}_{\infty}$ $\leq\mathrm{Lip}({\text{MLP}})_{\infty}\cdot\norm{\vec{e}_{n_{ctx}}\text{E}B+S-\vec{e}_{n_{ctx}}\text{E}B^{\prime}+S^{\prime}}_{\infty}$
$\leq\mathrm{Lip}({\text{MLP}})_{\infty}\cdot\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
D.5 Proof of theorem 4.5
First we will restate theorem 4.5 for convenience:
Theorem D.7.
If
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$
$<\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)/2,$
then the output of model $\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed for all inputs in $\mathbb{O}_{des}$. Moreover, we can use algorithm 1 to produce an upper bound $W$ for $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$.
Proof.
We first make use of theorem 4.4 to prove the first statement of the above theorem. Now, we just need to prove that the bound, $W$ in algorithm 1, is a valid bound on the lift $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$. By the monotonicity of lifting, we then have that $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}\leq\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
First note that, by lemma C.1, we have that
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq\mathrm{Lip}(\mathrm{Unembed})_{\infty}\cdot(\mathcal{W}(f^{\texttt{AttnH}};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$+\mathcal{W}(f^{\text{MLP}};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}+\mathcal{W}(f^{\mathbb{I}};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty})$
where
$f^{{\texttt{comp}}}(X,(B,S))={\texttt{comp}}\circ f^{enc}(X,(B,S))$
and
$f^{enc}(X,(B,S))=(X\cdot\text{E}\cdot\mathrm{diag}(B)+S).$
Then, we can use lemma 4.12 to get that $W^{\texttt{Attn}}$ is a valid bound on the worst-case deviation of $f^{\texttt{AttnH}}$. Then, we use lemma 4.14 to get that $\mathrm{Lip}(\text{MLP})_{\infty}\cdot\mathcal{W}(f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})$ is a valid bound on the worst-case deviation of $f^{\text{MLP}}$. Finally, we use lemma 4.15 to get the bound on $f^{enc}$. ∎
Appendix E Appendix for Permutation Invariance
Algorithm 6 Naive Algorithm to find ${\alpha}_{free}^{\min}$.
To find ${\alpha}_{free}^{\max}$, switch the $\min$ to a $\max$ in the objective.
  Input: $f^{enc}$ and model $\mathcal{M}$ Let $f^{enc}(\vec{e})=\vec{e}\cdot\text{E}\cdot\mathrm{diag}(B)+S$
  for $k\in(n_{fix},{n_{ctx}}]$ do
      Set
$\vec{\ell}^{\min}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\min_{t\in\texttt{FreeToks}}f^{enc}(\vec{e}_{q})\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot f^{enc}(\vec{e}_{t})^{T}$
  end for
  Return ${\alpha}_{free}^{\min}=\sum_{k}\exp\mathopen{}\mathclose{{}\left(\vec{\ell}^{\min}_{k}}\right)$
E.1 Proof of Attention Worst-Case Deviation
In this subsection, we provide a proof of lemma 5.3 and lemma 5.4. First, we prove the correcntess of the algorithms to find ${\alpha}_{free}^{\min}$ and ${\alpha}_{free}^{\max}$.
Lemma E.1 (Restatement of lemma 5.3).
Both algorithm 3 and algorithm 6 provide valid bounds on ${\alpha}_{free}^{\min}$ and ${\alpha}_{free}^{\max}$.
Proof.
We provide a proof for ${\alpha}_{free}^{\min}$; the proof for ${\alpha}_{free}^{\max}$ is analogous.
First note that the blowup and shift are fixed and thus $f^{enc}$ is only a function of the its input token: i.e. we can rewrite $f^{enc}(\vec{e}_{t},BS)$ as $f^{enc}(\vec{e})=\vec{e}\cdot E\cdot\mathrm{diag}(B)+S$ for fixed $B$ and $S$. Then, for the naive algorithm, algorithm 6, the correctness follows from the fact that the minimum possible logit value from each position can be calculated by enumerating over all the free tokens.
For the LP-based algorithm, algorithm 3, we first consider the LP as an integer linear program. Consider the permutation matrix $X$ with entries $x_{j,t}$, where $x_{j,t}=1$ if token $t$ is selected for position $j$ and $x_{j,t}=0$ otherwise. Then, the LP can be seen as finding the permutation matrix which minizes the objective function, ${\alpha}_{free}^{\min}$. So, the relaxed LP is a lower bound on the integer LP which is a lower bound on the true value of ${\alpha}_{free}^{\min}$. ∎
We now restate the lemma for the worst-case deviation of attention.
Lemma E.2.
Worst-case deviation of attention is bounded as follows for fixed blowup $B$ and shift $S$ when considering the permutation class $[X]$ with free tokens FreeToks:
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot B+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\norm{(E[(n_{fix},{n_{ctx}}),:]B+S)\cdot V}_{{F_{\infty}}}.$
Proof.
The proof follows analogously to the proof of lemma 4.12 (bound on worst-case deviation of attention in the generic case). The main difference is that we have a singleton set for the blowup and shift and that we restrict the free tokens. First, this implies that the restriction of $\mathcal{B},\mathcal{S}$ to singelton sets means that no optimization over blowup and shift are needed in the bounds as the blowup and shift become constants. Second, the restriction of the free tokens means that the worst-case contribution from the value function, $(E\mathrm{diag}(B)+S\cdot V)$, is restricted to the embeddings selected by the free tokens. So, we replace $\norm{E\mathrm{diag}(B)+S\cdot V}_{{F_{\infty}}}$ with $\norm{(E[(n_{fix},{n_{ctx}}),:]\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}$. Moreover, $\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S})_{\infty}=0$ as the blowup and shift are fixed and so we can drop the last term in the bound in lemma 4.12. ∎
E.2 Proof of theorem 5.2
We first restate the theorem for convenience.
Theorem E.3.
If
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};[X])_{\infty}<\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)/2$
then the output of model $\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed for all inputs in $[X]$. Moreover, we algorithm 2 in appendix E produces an upper bound $W$ for $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$.
Proof.
The proof follows analogously to the proof of theorem 4.5 in section D.5. The key difference is that, due to the fixing of the blowup and shift, the bound on $W$ simplifies. First, the worst-case deviation of $\vec{e}_{{n_{ctx}}}\cdot f^{enc}(X)$ is 00 as the blowup and shift are fixed. So, the worst-case deviation of the MLP and identity function are 00 as well. Then, algorithm uses either algorithm 6 or algorithm 3 to find bounds on the extremal values of attention with provable correctness as per lemma 5.3. Next, the algorithm simplified worst-case deviation of attention (as per lemma 5.4) to bound the worst-case deviation of attention. Finally, the algorithm uses the Lipschitz constant of the unembedding function, invoking lemma C.1, to bound the worst-case deviation of the model. ∎
Appendix F Overwhelming for ${n_{ctx}}\to\infty$
In this section, we will consider a specific set ${s}=\bigtimes_{n_{fix}}\vec{e}_{r}$ and $q=\vec{e}_{r}$ for some fixed $r\in[{d_{vocab}}]$. In words, we will consider the fixed input to be one repeated token. We term this a “repetition restriction.”
Moreover, it will be useful to define the set of all possible inputs under a repetition restriction.
Definition F.1 (Repetition Space).
Let $\mathbb{O}_{rep}\subset\mathbb{O}^{n_{ctx}}$ be the set of all possible inputs under a repetition restriction. That is,
$\mathbb{O}_{rep}=\mathopen{}\mathclose{{}\left\{X\in\mathbb{R}^{{n_{ctx}}\times{d_{vocab}}}\mid X=\begin{bmatrix}\vec{e}_{r}^{T}\\ \vec{e}_{r}^{T}\\ \vdots\\ \vec{e}_{r}^{T}\\ Y\\ \vec{e}_{q}\end{bmatrix},Y\in\mathbb{O}^{n_{free}}}\right\}.$
For simplicity, we will also not consider positional encodings in this section. I.e. we remove the use of RoPE in the attention mechanism. Though as Ref. (Barbero et al., 2024) pointed out, rotary positional encodings (Su et al., 2024) converge to providing zero information as ${n_{ctx}}\to\infty$.
Theorem F.2 (Asymptotic Convergence to a Fixed Model).
If $\frac{n_{free}}{{n_{ctx}}}\in o(1)$ as a function of ${n_{ctx}}$, then the repetition restriction converges to a fixed model if $\mathrm{PTP}(\mathcal{M},X)$ is positive for $X=\vec{e}_{r}$.
To prove this theorem, we will need to (a) find a way to compute a $\mathrm{PTP}$ function of the model and input independent of ${n_{ctx}}$ and (b) use the framework of section 4 to individually bound the worst-case deviation for each component of the model.
Computing Peak-to-Peak Difference
Given that $\mathrm{PTP}$ is computed by evaluating the model on a single input, finding a $\mathrm{PTP}$ value is normally a simple task. But, in the asymptotic case, we need to “shortcut” the computation of $\mathrm{PTP}$ for an $X\in\mathbb{O}_{rep}$. We can simply do this by setting all the free tokens to the repetition token: $X={\vec{e}_{r}}^{n_{free}}$.
Lemma F.3 (Shortcut for $\mathrm{PTP}$).
We can compute $\mathrm{PTP}(\mathcal{M},X)$ in time $O(1)$ for $X=\vec{e}_{r}^{{n_{ctx}}}$ even as ${n_{ctx}}\rightarrow\infty$.
Proof.
We can note that $[X]=\vec{e}_{r}\cdot\text{E}$ E and $\mathrm{Var}[X]=\mathrm{Var}[\vec{e}_{r}\cdot\text{E}]=0$. So, we can easily compute blowup, $B$, and shift, $S$. Note, for attention head $\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q}$, the output equals
$\vec{p}(X)\cdot(XE\cdot\mathrm{diag}(B)+S)\cdot V$
for some probability vector $\vec{p}(X)$. Because $X=\vec{e}_{r}^{n_{ctx}}$, $(XE\cdot\mathrm{diag}(B)+S)\cdot V=\vec{a}^{n_{ctx}}$ for vector $\vec{a}=(\vec{e}_{r}\cdot E\cdot\mathrm{diag}(B)+S)V$. Thus,
$\vec{p}(X)\cdot(XE\cdot\mathrm{diag}(B)+S)\cdot V=\vec{a}.$
We therefore get that the output of $\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed to $\vec{e}_{r}(\text{E}\cdot\mathrm{diag}(B)+S)\cdot V$ for all context windows ${n_{ctx}}$. Finally, we can simply compute the value of $\vec{e}_{n_{ctx}}\cdot f^{\text{MLP}}_{(n_{fix},{n_{ctx}})\mid{s},q}$ because we know $B$ and $S$ and so just need to compute $\text{MLP}(\vec{e}_{r}(E\cdot\mathrm{diag}(B)+S))$. ∎
Bounds on Expectation and Variance
Lemma F.4 (Bounds on Expectation and Variance).
Let $X^{\prime}\in\mathbb{O}_{rep}$ and $X=\mathopen{}\mathclose{{}\left(\bigtimes_{i=1}^{n_{fix}}\vec{e}_{r}}\right)\times X^{\prime}\times\vec{e}_{q}$. I.e. $X$ is an expansion of the input with the first tokens being the repetition token and the last being the query token. Then, we can bound the column wise expectation for column $j$, $\mu_{j}$, and variance, $\mathrm{Var}_{j}$ of $X$ by
$\frac{1}{{n_{ctx}}}\mathopen{}\mathclose{{}\left(n_{fix}\cdot\vec{e}_{r}\text{E}\vec{e}_{j}^{T}+n_{free}\min_{i}(\vec{e}_{i}\text{E}\vec{e}_{j}^{T})+\vec{e}_{q}\text{E}\vec{e}_{j}^{T}}\right)\leq\mu_{j}\leq$
$\frac{1}{{n_{ctx}}}\mathopen{}\mathclose{{}\left(n_{fix}\cdot\vec{e}_{r}\text{E}\vec{e}_{j}^{T}+n_{free}\max_{i}(\vec{e}_{i}\text{E}\vec{e}_{j}^{T})+\vec{e}_{q}\text{E}\vec{e}_{j}^{T}}\right).$
Let $\mu_{j}^{\min}$ and $\mu_{j}^{\max}$ denote the lower and upper bounds. Then, we can bound the variance of the $j$-th column of $\vec{x}\text{E}$ by
$\mathrm{Var}_{j}\leq$ $\frac{1}{{n_{ctx}}}\cdot\max_{\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]}\bigg{[}(n_{fix}+1)\cdot\mathopen{}\mathclose{{}\left(\vec{e}_{r}\text{E}\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+n_{free}\cdot\max_{i\in{d_{vocab}}}\mathopen{}\mathclose{{}\left(\vec{e}_{i}\text{E}\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}\bigg{]}.$
We define $\mathrm{Var}_{j}^{\min},\mathrm{Var}_{j}^{\max}$ to be the lower and upper bounds respectively.
Proof.
The bounds on $\mu_{j}$ follow as we minimize and maximize the contribution of each free row to the expectation of the column of $j$. The proof of for variance bounds follows similarly. For the lower bound, we note that the variance contribution from each row is at least 00. For the upper bound, we simply maximize the contribution of each row to the variance. ∎
Recall the definition of blowup and shift sets from definition 4.6.
Lemma F.5 (Blowup and Shift Set Bounds).
Recall that $\gamma$ and $\beta$ are the learned constants for layer normalization. Let $\vec{x}\in\mathbb{O}_{des}$. Then, we can bound the blowup and shift sets for the $j$-th column of $\vec{x}\text{E}$. For all $S_{j}\in\mathcal{S}_{j}$ and $B_{j}\in\mathcal{B}_{j}$, we have
$\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\max}+\epsilon}}\leq B_{j}\leq\frac{\gamma}{\sqrt{\epsilon}}$
Let $B_{j}^{\min}$ and $B_{j}^{\max}$ denote the lower and upper bounds respectively. Then,
$B_{j}^{\min}\mu_{j}^{\min}+\beta\leq S_{j}\leq B_{j}^{\max}\mu_{j}^{\max}+\beta.$
Now as a corollary of lemma F.4 and lemma F.5, we can show that the minimum and maximums of the blowup and shift sets can be made arbitrarily close together.
Corollary F.6 (Convergence of Blowup and Shift Sets).
If $\frac{n_{free}}{{n_{ctx}}}\in o(1)$, then for every $\delta_{1},\delta_{2}\in(0,1)$, there exists some setting of ${n_{ctx}}^{\prime}$ such that for all ${n_{ctx}}>{n_{ctx}}^{\prime}$,
$B_{j}^{\max}-B_{j}^{\min}\leq\delta_{1}\quad\text{and}\quad S_{j}^{\max}-S_{j}^{\min}\leq\delta_{2}.$
Proof.
The corollary follows from the fact that $\mu_{j}^{\min},\mu_{j}^{\max}$ converge to $\vec{e}_{r}E\vec{e}_{j}^{T}$ as ${n_{ctx}}\rightarrow\infty$. So then, $(\vec{e}_{r}E\vec{e}_{j}^{T}-\mu^{\prime})^{2}$ for all $\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]$ converges to 00 and thus the upper bound on $\mathrm{Var}_{j}$ converges to 00 as long as $\frac{n_{free}}{{n_{ctx}}}$ converges to 00. ∎
Therefore the blowup and shift sets converge, let $B_{j}^{\prime}$ and $S_{j}^{\prime}$ denote the converged values for $B_{j}$ and $S_{j}$ respectively.
Now, we need to show a bound on the difference between the maximum logit for the free tokens and the minimum logit for the repeated tokens in the attention head.
Bound on Attention Head Difference
To bound attention, we will take advantage of the repeated structure as well as the ideas in lemma 4.11 (bounds on the attention weights).
Corollary F.7 (Sum of Attention Weights).
Let $L_{fix}^{\min}\leq\min_{i\in[n_{fix}]\cup\{n_{c}tx\}}\ell_{i}$ and $L_{free}^{\min}\leq\min_{j\in(n_{fix},{n_{ctx}})}\ell_{j}$. Also, let $L_{fix}^{\max}\geq\max_{i\in[n_{fix}]\cup\{n_{c}tx\}}\ell_{i}$ and $L_{free}^{\max}\geq\max_{i\in(n_{fix},{n_{ctx}})}\ell_{i}$. Then,
$\frac{n_{fix}}{n_{fix}+1+(n_{free})\cdot e^{L_{free}^{\max}-L_{fix}^{\min}}}\leq\sum_{i\in[n_{fix}]\cup\{{n_{ctx}}\}}\text{softmax}(\vec{\ell}_{i})\leq\frac{n_{fix}}{n_{fix}+1+(n_{free})\cdot e^{L_{free}^{\min}-L_{fix}^{\max}}}$
and
$\frac{n_{free}}{n_{fix}+1+(n_{free})\cdot e^{L_{free}^{\min}-L_{fix}^{\max}}}\leq\sum_{j\in(n_{fix},{n_{ctx}})}\text{softmax}(\vec{\ell}_{j})\leq\frac{n_{free}}{n_{fix}+1+(n_{free})\cdot e^{L_{free}^{\max}-L_{fix}^{\min}}}.$
Lemma F.8 (Bound on Difference).
We can bound the difference between the maximum logit for the free tokens and the minimum logit for $n_{fix}$ repeated tokens
$L_{free}^{\max}-L_{fix}^{\min}$ $\leq\theta$ (5)
and
$L_{free}^{\min}-L_{fix}^{\max}$ $\leq\theta$ (6)
where $\theta$ is a small constant.
Proof.
We will prove the first inequality as the second follows analogously. Assume that $i\in[{d_{vocab}}]$ is the token which maximizes the difference. We will relly on the convergence of the blowup and shift sets to show that the difference between the maximum logit for the free tokens and the minimum logit for the repeated tokens in the attention head is bounded.
Connote $\mathcal{B}-B^{\prime}$ as the set of blowup values shifted by the converged value and $\mathcal{S}-S^{\prime}$ as the set of shift values shifted by the converged value. Then $\mathcal{B}-B^{\prime}$ and $\mathcal{S}-S^{\prime}$ are bounded by $\delta_{1},\delta_{2}$ respectively, we can write the following bound on the difference.
$L_{free}^{\max}-L_{fix}^{\min}$ $\leq\max_{\epsilon\in\mathcal{B}-B^{\prime},\;\epsilon_{S}\in\mathcal{S}-S^{\prime}}(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime}+\epsilon)$
$\cdot QK^{T}(\mathrm{diag}(B^{\prime}+\epsilon)\cdot E^{T}\vec{e}_{q}^{T}+S^{T}+\epsilon_{S})$
$=\max_{\epsilon\in\mathcal{B}-B^{\prime},\epsilon_{S}\in\mathcal{S}-S^{\prime}}$
$\sum_{d_{1}\in\{\epsilon,B^{\prime}\}}\sum_{d_{2}\in\{\epsilon,B^{\prime}\}}\sum_{d_{3}\in\{\epsilon,S^{\prime}\}}(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(d_{1})\cdot QK^{T}(\mathrm{diag}(d_{2})\cdot E^{T}\vec{e}_{q}^{T}+d_{3})$
$\leq\max_{\epsilon\in\mathcal{B}-B^{\prime},\epsilon_{S}\in\mathcal{S}-S^{\prime}}(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\epsilon,\epsilon_{S})$
$\leq(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\delta_{1},\delta_{2},B^{\prime},S^{\prime})$ (By the bounds on the blowup and shift set)
$\leq(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\delta_{1},\delta_{2})$
where the last inequality follows from considering $B^{\prime}$ and $S^{\prime}$ as a constant. Then, because $\delta_{1},\delta_{2}$ decrease as a function of ${n_{ctx}}$ and corollary F.7, we can bound
$(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\delta_{1},\delta_{2})\leq\theta$
for a large enough ${n_{ctx}}$. ∎
Lemma F.9 (Attention Bound, lemma 4.12).
For large enough ${n_{ctx}}$, we get
$\mathcal{W}(f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})_{\infty}\leq o(1).$
Proof.
Recall that lemma 4.12 gives us
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\max_{B,S}\norm{\vec{e}_{r}\text{E}B+S}+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(\vec{e}_{r}\text{E}B+S)\cdot V}_{\infty}$
Because $L_{free}^{\max}-L_{fix}^{\min}\leq\theta$ and $L_{free}^{\min}-L_{fix}^{\max}\leq\theta$, we can use corollary F.7 to upper-bound ${\beta}_{fix}^{\max}-{\beta}_{fix}^{\min}$ and ${\beta}_{free}^{\max}$ by
$1-\frac{n_{fix}}{n_{fix}+(n_{free}+1)\cdot\theta}=\frac{(n_{free}+1)\cdot\theta}{n_{fix}+(n_{free}+1)\cdot\theta}\leq\frac{n_{free}\cdot\theta}{n_{fix}}\leq\theta\cdot o(1).$
So then
$\mathcal{W}(f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq o(1)$
as $\theta\in o(1)$. ∎
Bound on $f^{enc}$
Finally, as per section 4, we need to bound the worst-case deviation of the $f^{enc}$ function. Given that $f^{enc}$ is a linear operation as a function of blowup and shift sets, we can use the convergence of the blowup and shift sets to show that the worst-case deviation of $f^{enc}$ is bounded.
Lemma F.10 (Bound on $f^{enc}$).
For large enough ${n_{ctx}}$, we have
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})_{\infty}\leq{\mathsf{poly}}(\delta_{1},\delta_{2}).$
, where $\delta_{1},\delta_{2}$ are the bounds on the blowup and shift differences.
We are now ready to prove theorem F.2.
Proof of theorem F.2.
Note that we reduced $\mathcal{W}(\vec{e}_{n_{ctx}}\cdot\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})_{\infty}$ to be upper-bounded by $O(1)\cdot\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})+o(1)$ through lemma 4.14 and lemma F.9. Then, by lemma F.10, we have that theorem F.2 holds as $\mathcal{W}$ goes to 00 as ${n_{ctx}}\rightarrow\infty$. So, as long as $\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)$ is positive for some $X\in\mathbb{O}_{rep}$, then we converge to “overwhelming” by theorem 4.4. Note that $\vec{e}_{r}^{n_{ctx}}\in\mathbb{O}_{rep}$ and by lemma F.3, we can compute a sample of peak-to-peak deviation for all ${n_{ctx}}$. ∎